November 30, 2005

Status report on Star Tracker and TTCS (Corrado Gargiulo)
Corrado informs that the star tracker support complete of lens and electronics underwent to vibration test at Terni CERMIS facility. The test was successful and the first vibration frequency is located at 60 Hz, above the threshold of 50 Hz.

The TTCS box (TTCBP) layout is terminated and the Chinese group is performing the FEA calculations.

Drawings at

Status report on Gas Circuit Control (Francesca Bucci, Francesca Spada)
Bruno reports that last Tuesday Alessandro has looked with scope the noise on pressure sensor P1. It’s evident a fast signal of 300 kHz repetition rate, of unknown origin and about 0.3 V. This noise is accompanied by a 50Hz power noise of about 100 mV. Clearly the 50 Hz noise is affected by changing wires routing. Signal wires from pressure sensors will be wired in twisted fashion.


F. Bucci - "MCA Status"
F. Spada "Gas System Status"
M. Sapinski "Cuts on momentum reconstruction for antideuteron analysis"